# This is a preprint submitted to: Computer Methods in Applied Mechanics and Engineering

### v1

@@ 5 Discussion
One of the clear benefits of using extended versions of PME is the ability to incorporate lumped physical metrics and distributed fields, enabling a tighter link between geometry and performance. All four test cases define lumped metrics, each representing integral performance indicators at specific operating conditions. By embedding these scalars, low-fidelity predictions of these lumped metrics can be tightly coupled with the reduced design variables from PME-based methods, simplifying multi-fidelity optimization. Furthermore, early sensitivity analyses are made easier by looking at, e.g., the 2D contour plots (see Fig. 13) reveal how varying just one or two reduced coordinates shifts integral performance (e.g., drag, thrust, seakeeping motion) away from the baseline configuration. This “lumped-level” lens is invaluable for quick screening, where high-fidelity CFD simulations are too expensive to run exhaustively. By building on these low-fidelity models, designers can prune unpromising shape variations early on. Furthermore, this low-fidelity model can be already used to illustrate the trade-offs between multiple objectives (e.g., lowering drag vs. improving lift, increasing thrust vs. reducing vortex intensity, or minimizing calm-water resistance while controlling seakeeping motions), as well as providing a first Pareto front of optimal solutions (see Fig. 14).
@@
Alongside lumped metrics, the embedding often includes distributed fields—pressure over the surface for the airfoil and glider, or near-blade flows for the propeller. These data prove critical if the final goal is to train machine-learning reduced-order models (ROMs) capable of predicting integral quantities of interest. For instance, for the RAE-2822, the pressure distributions can be aggregated into a small set of modes that reconstruct aerodynamic loads, while for the propeller, it provides richer mode shapes and ensures variations that affect vortex shedding, cavitation risk, or tip flow details. Since the embeddings reduce the dimension of these distributed fields, surrogate ROMs can be trained more efficiently, predicting crucial integral outputs without re-running the solver for every shape perturbation.

### v2

@@ 5 Discussion
One of the primary motivations behind the proposed PI-/PD-PME extensions is to incorporate physical insight into the dimensionality reduction process without imposing prohibitive computational costs. While enriching the embedding with physical quantities (e.g., lift, drag, or pressure fields) might suggest an increased overhead, these quantities are obtained from low-fidelity solvers (e.g., XFOIL and PUFFIn), which are computationally inexpensive and widely used in early-stage design. Importantly, in many practical workflows, these evaluations are already available from previous parametric studies, and thus no new high-fidelity simulations are required for training. This enables a cost-effective embedding that reflects relevant physical trends.
This rationale is further supported by an empirical analysis of variance convergence. For all test cases considered, the evolution of both geometric and physical variance was monitored as a function of the number of Sobol samples used in the training dataset. Despite generating up to 16,384 samples, it was observed that variance saturation typically occurs around 1,000 samples. This indicates that the main informative directions in the design space can be captured with a limited number of low-fidelity simulations. Such sampling budgets are compatible with standard preliminary design campaigns, even for problems involving more than 20 design variables. The convergence behavior is illustrated in Fig. LABEL:fig:var_conv, where both geometric and physical variance show rapid stabilization across all configurations. Once this saturation is reached, the structure of the reduced manifold becomes largely insensitive to additional samples, supporting the robustness of the method.
Moreover, investing in a physics-informed latent space leads to significant downstream savings. By embedding performance-related directions early, the method avoids wasting computational effort on uninformative or non-promising shape perturbations. This is confirmed, for instance, in the propeller case study, where the dominant reduced basis direction consistently aligns with pitch at tip, a shape feature known to have strong influence on hydrodynamic performance. Similar trends are observed across other configurations, indicating that the reduced space prioritizes design variations with high physical relevance. This is particularly valuable when transitioning to high-fidelity or multi-objective optimization, where the cost of full evaluations is orders of magnitude higher. Unlike purely geometric PCA or POD, which may generate reduced spaces uncorrelated with performance, or black-box regressors that do not provide a true reduction of the design space, PI-/PD-PME yield interpretable, physics-aware coordinates that guide the design exploration more effectively.
@@
It is also important to note that the proposed dimensionality reduction does not require highly accurate physical outputs: it relies on capturing global variance trends rather than absolute values. As long as the low-fidelity solver provides coherent gradients and relative sensitivities across the design space, the method effectively identifies the most informative directions. This assumption is supported by common practice in design optimization, where low-fidelity models are often benchmarked against high-fidelity responses along selected design directions to validate their qualitative behavior. Moreover, the dimensionality of the reduced space is deliberately kept higher than strictly necessary to mitigate potential inaccuracies in the low-fidelity-based embedding and to retain sufficient flexibility for high-fidelity refinement. This allows the embedding to serve not only as a compact representation of the space, but also as a candidate low-fidelity model in a multi-fidelity optimization framework. Furthermore, early sensitivity analyses are made easier by looking at, e.g., the 2D contour plots (see Fig. LABEL:fig:contours), which reveal how varying just one or two reduced coordinates shifts integral performance (e.g., drag, thrust, seakeeping motion) away from the baseline configuration. This “lumped-level” lens is invaluable for quick screening, where high-fidelity CFD simulations are too expensive to run exhaustively. By building on these low-fidelity models, designers can prune unpromising shape variations early on. Furthermore, this low-fidelity model can be already used to illustrate the trade-offs between multiple objectives (e.g., lowering drag vs. improving lift, increasing thrust vs. reducing vortex intensity, or minimizing calm-water resistance while controlling seakeeping motions), as well as providing a first Pareto front of optimal solutions (see Fig. LABEL:fig:paretos).
@@
Alongside lumped metrics, the embedding often includes distributed fields—pressure over the surface for the airfoil and glider, or near-blade flows for the propeller. These data prove critical if the final goal is to train machine-learning reduced-order models (ROMs) capable of predicting integral and/or local quantities of interest. For instance, the pressure distribution, for the RAE-2822, can be aggregated into a small set of modes that reconstruct aerodynamic loads, while, for the propeller, it provides richer mode shapes and ensures variations that affect vortex shedding, cavitation risk, or tip flow details. Since the embeddings reduce the dimension of these distributed fields, surrogate ROMs can be trained more efficiently, predicting crucial integral outputs without re-running the solver for every shape perturbation.
Finally, a distinctive feature of the PI-/PD-PME framework is its ability to accommodate a large number of physical quantities without requiring prior manual selection. When the most informative observables are not known in advance, the method identifies dominant contributors based on their empirical variance and statistical correlations. It should be noted, however, that physical metrics that are redundant or irrelevant for performance are not automatically downweighted: if repeated or exhibiting high variance, they may still dominate the embedding. As such, proper normalization and careful feature selection remain important to avoid unintentional overemphasis. Nonetheless, this unsupervised, data-driven formulation enables the identification of performance-relevant directions and supports the exploration of complex design spaces where physical intuition alone may be insufficient. The relative impact of each physical feature on the reduced modes is visualized in Figs. 1114.
@@ 6 Conclusions
Finally, while the proposed methodology effectively incorporates physical information into the reduced space, a limitation lies in its inherently linear structure. In practical applications, both geometric deformations and physical responses often exhibit nonlinear behavior that may not be fully captured by linear embeddings. PME does not impose explicit assumptions regarding the smoothness or continuity of the geometric or physical fields. Its basis construction is entirely data-driven and reflects the empirical variance across samples. As a linear technique, PME can still capture sharp gradients or localized features—provided they manifest as dominant directions of variability. However, in design spaces characterized by strong discontinuities or highly nonlinear relationships, its effectiveness may be limited, and nonlinear extensions could provide more expressive representations. Therefore, future research should explore nonlinear extensions of PME by leveraging advanced dimensionality reduction techniques, such as local PCA [54], kernel PCA [31, 55], autoencoders [56, 57, 58, 59, 60, 61], or principal geodesic analysis [62, 63], which allow for the discovery of curved low-dimensional manifolds. These techniques may further enhance the ability of the framework to capture complex dependencies between design variables and physical observables, improving both the fidelity and generalizability of the reduced representations. Nonetheless, to the best of the authors’ knowledge, none of these nonlinear approaches currently support the core capabilities offered by PME—namely, the ability to assign custom weights across heterogeneous features and to perform analytical backmapping to the original parametric space. Developing nonlinear extensions of PI-/PD-PME that retain such properties represents a compelling and largely unexplored direction for future research.